OpenStudy (nesha97):

Write a function with the following characteristics 20. A vertical asymptote at x= -1 An oblique asymptote at y = x + 2

OpenStudy (tkhunny):

You should not be struggling with the first. Please demonstrate that one and we can talk about the second. ALWAYS show YOUR work int he VERY FIRST post on a thread.

OpenStudy (nesha97):

well i know for the vertical asymptote that x+1 would be the denominator right???

OpenStudy (faiqraees):

Correct

OpenStudy (faiqraees):

For oblique asymptote, the general method is \[f(x) = oblique~asymptote +\frac{numerator}{horizontal~asymptote} \] where numerator has to be an expression with a variable degree of less than that of numerator

OpenStudy (nesha97):

so what do i put for the horizontal asymptote and the numerator?

OpenStudy (faiqraees):

You can put any constant for a numerator For horizontal asymptote, you have to put x+1

OpenStudy (nesha97):

x+2+ 2/x-1 like this??

OpenStudy (tkhunny):

No, but only because you have not observed properly the Order of Operations.

OpenStudy (faiqraees):

\[f(x)=x+2+\frac{2}{x+1}\] would be correct

OpenStudy (nesha97):

so thats it?? dont i have to write it back into a rational function??

OpenStudy (faiqraees):

Yeah you can do that too

OpenStudy (nesha97):

how would i do that

OpenStudy (faiqraees):

\[y = \frac{(x+1)(x+2) +2}{x+1} \]

OpenStudy (nesha97):

Ok i factored the numerator and added the 2 and got x^2+3x+4 for the numerator and x+1 for the denominator.....Thank you so much I've been stressed out with that problem since forever